find all zeros of the function. f(x)=x^3-3x^2-6x+8
Answer by ewatrrr(24785)


Hi
finding all zeros of the function.
f(x)=x^3-3x^2-6x+8
x^3-3x^2-6x+8 = 0 |If x = 1 then zero (1-3-6+8 = 0 )
(x -1)(x-4)(x+2) =0 |Dividing x^3-3x^2-6x+8 by (x-1)
( x-1)=0 x = 1
(x-4)=0 x = 4
(x+2)=0 x = -2